# Evaluating the Utility of Conformal Prediction Sets for AI-Advised Image Labeling

### v1

@@ Abstract.
As deep neural networks are more commonly deployed in high-stakes domains, their lack of interpretability makes uncertainty quantification challenging. We investigate the effects of presenting conformal prediction sets—a method for generating valid confidence sets in distribution-free uncertainty quantification—to express uncertainty in AI-advised decision-making. Through a large pre-registered experiment, we compare the utility of conformal prediction sets to displays of Top-1 and Top-k predictions for AI-advised image labeling. We find that the utility of prediction sets for accuracy varies with the difficulty of the task: while they result in accuracy on par with or less than Top-1 and Top-k displays for easy images, prediction sets excel at assisting humans in labeling out-of-distribution (OOD) images especially when the set size is small. Our results empirically pinpoint the practical challenges of conformal prediction sets and provide implications on how to incorporate them for real-world decision-making.
@@ 1. Introduction
As machine learning models like deep Neural Networks (NNs) achieve impressive predictive performance, AI-advised decision-making has become more commonplace in various tasks where a human must make a decision, from high stakes domains like medicine (e.g., (Razzak et al., 2018)) or autonomous vehicle navigation (e.g., (Fujiyoshi et al., 2019)) to online tasks like content moderation (e.g., (Lai et al., 2022)) or image labeling (e.g., (van der Wal et al., 2021)). Ideally, having access to a predictive model can help humans to make more informed decisions, surpassing the capabilities of either human judgment or AI in isolation, a synergy known as complementary performance (Bansal et al., 2021; Hemmer et al., 2021). Whenever an AI system understands certain regions of the feature space better, the human counterpart can benefit by considering the model predictions. For example, in AI-advised image labeling, which we study in this paper, human labelers may benefit from viewing predictions from a model trained on previously labeled examples.
@@
To help humans know when to trust model predictions, presenting information about the probability that the model’s prediction is correct can in principle be useful, such as providing the softmax pseudo-probability produced in the final layer of NNs for classification tasks. However, predictions from NNs are known to be overconfident (Guo et al., 2017), and heuristic measures of uncertainty are poorly calibrated. Consequently, decision-making relying on definitive NN predictions (e.g., Top-) with the highest softmax prediction confidence can lead to an unwarranted impression that prediction uncertainty is minimal or not important, affecting the quality of the decisions that are ultimately made.
@@
Given these challenges, conformal prediction (Vovk et al., 2005; Romano et al., 2020; Angelopoulos et al., 2020) has emerged as an alternative solution that can rigorously quantify the prediction uncertainty for NNs through the use of prediction sets. Instead of presenting unreliable softmax values or other heuristics for uncertainty, the prediction set is a type of confidence interval, consisting of a set of predictions with a coverage guarantee—the true class is, on average, captured within such sets with a user-specified probability (e.g., ). The conformal framework is compatible with NNs because the derived sets are distribution-free—they are model-agnostic, work with finite samples, and offer non-asymptotic guarantees without distributional assumptions (Angelopoulos and Bates, 2021).
@@
One challenge is that while conformal prediction sets express prediction uncertainty with guarantees that status quo presentations like Top- predictions with softmax values cannot, the conformal coverage guarantee can lead to large sets embodying high uncertainty for instances the model perceives as difficult (Babbar et al., 2022; Angelopoulos et al., 2020). Compared with Top- predictions, such sets may increase cognitive load, rendering uncertainty quantification less effective (Zhou et al., 2017), and thus compromising predictive performance. To understand how conformal prediction sets support human decisions aided by model predictions in an image labeling task, we contribute a large pre-registered online experiment that compares the accuracy achieved in AI-advised image labeling using conformal prediction sets with that achieved with no predictions or Top- predictions. Participants in our study were tasked with labeling images from ILSVRC 2012 (Deng et al., 2009) that varied in difficulty, and encompassed both image stimuli that were in-distribution and out-of-distribution (OOD). This allowed us to evaluate the utility of prediction sets against Top- presentation alternatives in scenarios where the model’s predictions are produced by a well-calibrated model and when predictions are unreliable (i.e., more error-prone), such as when models are poorly calibrated due to shifts in the distribution of new instances relative to training data.
@@
We evaluate decision-making by metrics such as accuracy and the shortest path length in the label space hierarchy, which approximates the amount of deviation from the ground truth in the label space hierarchy. We also use participants’ elicited willingness-to-pay at the end of the experiment to compare the perceived value of predictions against the post-hoc inferred monetary benefits derived from accessing these predictions.
@@
We find that for in-distribution instances, prediction sets lead to reduced labeling accuracy compared to Top- predictions. However, for OOD instances where the model misspecification differentially affected the calibration of the Top- predictions, prediction sets improve accuracy regardless of set size. We find that participants are willing to pay roughly equivalent amounts for each type of display; if anything, they undervalue prediction sets relative to other presentations despite their performance advantages. Our results advance understanding of the strengths and limitations of prediction sets as a more robust form of uncertainty quantification for improving AI-advised decisions, and highlight how decision-makers can over-rely on provided predictions, even when poorly calibrated, in an image labeling task.
@@ 2.1. Uncertainty Quantification Using Conformal Prediction
Prior work on Human-Computer Interaction (HCI) suggests that humans can make more informed decisions when uncertainty is effectively communicated (e.g., (Joslyn and LeClerc, 2012; Kay et al., 2016; Hullman et al., 2018)), such as by presenting prediction uncertainty through static intervals (Cumming, 2014; Cumming and Finch, 2005; Manski, 2019; Taylor and Kuyatt, 1994) or animations (Hullman et al., 2015; Zhang et al., 2021). However, uncertainty quantification for NNs remains challenging (Amodei et al., 2016; Angelopoulos and Bates, 2021) due to their poor interpretability and often unreliable softmax pseudo-probabilities (Guo et al., 2017; Torralba and Efros, 2011).
@@
One approach to quantify prediction uncertainty in classification tasks is to construct Bayesian NNs, so that predictions, which depend on sampling from the posterior distributions of model weights, carry an inherent, mathematically-grounded measure of uncertainty. However, Bayesian NNs heavily rely on distributional assumptions for their priors and can be computationally intensive to train. Conformal prediction (Vovk et al., 2005) has emerged as a more flexible and efficient solution that does not require a strong model or distributional assumptions. In classification settings, conformal prediction quantifies uncertainty through prediction sets. Similar to traditional confidence intervals, the derived set contains predicted classes and achieves a coverage guarantee that, on average, the true class is captured by the set with a user-specified probability. Formally, given a training set , where represents the feature vector and represents its corresponding label, the uncertainty set function, denoted , maps a feature vector to a subset of to ensure the coverage guarantee with an error rate of .
@@
If an uncertainty quantification approach is well calibrated, we naturally expect it to reflect greater uncertainty on instances that are more difficult for the model (e.g., OOD instances). With prediction sets (Romano et al., 2020; Angelopoulos et al., 2020), this manifests as adaptiveness: conformal prediction set will be larger when the task is more difficult for the model. Adaptiveness is realized through the inductive conformal method, which quantifies the prediction uncertainty of a chosen classifier without changing the model using a designated calibration set (Papadopoulos et al., 2002), in which instances are independently and identically distributed (i.i.d.) as those on which the model is trained.
@@
For each calibration instance, the predicted softmax scores are adjusted by Platt scaling (Platt, 1999; Guo et al., 2017; Nixon et al., 2019) for better interpretability and further regularized to penalize unlikely classes (Angelopoulos et al., 2020). These calibrated probabilities are then ranked in descending order, reflecting the classifier’s confidence in each class. A conformity score is computed by summing these ranked probabilities up to the true class, capturing the model’s accumulated confidence for the instance leading up to the correct classification. From the distribution of conformity scores of all calibration instances, a threshold is computed by taking the quantile, after a finite sample correction for robustness. This calibration threshold makes the prediction set adaptive and maintains the coverage guarantee. For a new and unknown instance, this calibration threshold decides the size of the prediction set by including all labels whose softmax scores exceed this calibration threshold. In our study, we opted for the Regularized Adaptive Prediction Sets (RAPS) algorithm (Angelopoulos et al., 2020) to ensure that the prediction sets of our experiment are as small as possible while achieving, on average, the desired coverage rate.
@@ 2.2. Effects of Presenting Uncertainty in AI-advised Decisions
Pairing humans with an AI system in complex decision-making scenarios can elevate performance beyond what either can achieve individually, especially when their strengths are complementary (Bansal et al., 2021; Steyvers et al., 2022). Particularly, in AI-advised image labeling, a plethora of tools and software exists for boosting human annotators’ speed and accuracy (Sager et al., 2021). However, performance can be influenced by how predictions are presented—specifically, whether uncertainty is effectively communicated or not. When uncertainty is well communicated, predictions appear to be more reliable and trustworthy to humans (Zhang et al., 2020; Lai and Tan, 2019; Beller et al., 2013). Knowing a model’s prediction accuracy or the data-generating process can make predictions more transparent and explainable, which encourages analytical thinking and reduces over-reliance on predictions (Zhang et al., 2020; Prabhudesai et al., 2023). However, uncertainty quantification does not guarantee improved decision-making. Cognitive biases, such as tendencies to under- or over-react to information relative to rational standards play a major role in how one interprets uncertainty (e.g., (Ambuehl and Li, 2018; Tversky and Kahneman, 1971)). Additionally, Zhou et al. (2017) suggest that uncertainty quantification can improve trust under low cognitive load but has been found to diminish trust when uncertainty is presented ambiguously, demanding a high cognitive load.
@@
There is limited empirical evidence on how prediction sets influence trust and accuracy in AI-advised decision-making. One exception is Babbar et al. (2022), who investigated the perceived utility and performance of humans using RAPS versus Top- predictions for labeling images from CIFAR-100 (Krizhevsky, 2009). They found that participants using RAPS reported higher trust and perceived utility than those using Top- predictions. However, they found a negligible difference in performance between the two groups, which may be due to their small study size (i.e., 15 in each group) and a lack of variation in their image stimuli. They suggested that smaller sets might lead to better accuracy, but this observation was based on a comparison between sets generated by two different conformal methods. Recent work by Straitouri and Rodriguez (2023) compared labeling performance between two decision support systems: one that only allowed participants to select a label value from the prediction set and another that allowed them to choose freely one of 16 labels. Their results suggest that when the model is well calibrated, decision-making solely based on prediction sets tends to be more reliable and accurate because of the coverage guarantee.
@@
In contrast to Babbar et al. (2022), we do a much larger experiment that evaluates the utility of different uncertainty prediction displays for in-distribution and OOD stimuli, each of which is systematically categorized by difficulty and set size of the associated conformal prediction set. This enables us to explore changes in performance by instance type. Instead of restricting agency as in Straitouri and Rodriguez (2023), we create a more realistic labeling scenario where participants can freely search for their preferred choice from a vast label space hierarchy, as inaccurate predictions may still be useful serving as clues for participants to identify the correct answer.
@@ 3.1. Overview
We conducted a large pre-registered111Pre-registration: https://aspredicted.org/6gh27.pdf mixed-design repeated measure experiment on Prolific to evaluate how different displays of prediction uncertainty impact performance on an AI-advised labeling task, including relative to performance without access to predictions. Figure 1 provides an overview of our stimuli generation process, which systematically varied whether the image instances were in-distribution or OOD and easy or difficult, as well as varying the size of the prediction sets produced by conformal prediction.
Experiment diagram demonstrating the pipeline and manipulations.
@@
Each participant in our experiment labels a set of 16 images, sampled from the ILSVRC 2012. We assign participants to one of four conditions representing different prediction displays: (1) no predictions (baseline), (2) Top- prediction with softmax, (3) Top- predictions with softmax, or (4) a prediction set (RAPS) (Angelopoulos et al., 2020).
@@
In real-world scenarios, the deployed model can encounter unexpected OOD instances where the distribution of input features differs from what was trained and calibrated. We vary whether images are in-distribution or OOD within subjects. Among the 16 stimuli, 4 are in distribution, while 12 are not, as we expect greater variation in participants’ accuracy for OOD.
@@
In our experiment, OOD instances notably lower the model’s Top- prediction accuracy. While in some scenarios conformal coverage guarantees can be compromised by OOD stimuli, for the corrupted images from the ILSVRC 2012 dataset that we use as OOD instances, the adaptiveness of RAPS leads to good coverage despite OOD. However, the size of the prediction sets in number of labels included becomes much larger when the instance is harder for the model. To evaluate the impact of set size on the effectiveness of prediction sets, we balance set size within subject by categorizing images as small or large based on the size of the prediction set generated for the image. This factor enables us to examine the performance of the prediction set between size levels, as previous work suggests that smaller set size is more preferable (Romano et al., 2020; Angelopoulos et al., 2020) and may yield higher accuracy for in-distribution images (Babbar et al., 2022). We expect small set sizes to be more useful in general, as they reduce cognitive load while maintaining the conformal coverage guarantee for in-distribution images. Large sets embodying high uncertainty may still be useful for labeling when participants are highly uncertain of what the image shows and consequently heavily reliant on prediction sets. In our setting, set size can also be viewed as a nuanced measure of difficulty for in-distribution stimuli because images that require a larger set size to achieve the desired coverage are intrinsically more difficult (Angelopoulos et al., 2020).
@@
For both in-distribution and OOD stimuli, we balance difficulty within subject by easy and hard, using the median cross-entropy loss of in-distribution images as the threshold. The former enables us to assess whether participants are more inclined to accept predictions when they express high confidence. The latter allows us to examine whether participants are more likely to modify or reject low-confidence predictions and instead rely on their own intuition.
@@
Our study evaluates participants’ performance by their labeling accuracy and the shortest path length between the chosen label and the correct label in the WordNet hierarchy (Miller, 1995; Fellbaum, 1998). After they labeled all images, we asked them to report their willingness-to-pay for the prediction display they used (a value ranging from to USD, the maximum they could earn) if they were to repeat the study with a similar distribution of instances. Finally, we elicited the strategies employed to approach tasks.

### v6

@@ Abstract.
As deep neural networks are more commonly deployed in high-stakes domains, their black-box nature makes uncertainty quantification challenging. We investigate the presentation of conformal prediction sets—a distribution-free class of methods for generating prediction sets with specified coverage—to express uncertainty in AI-advised decision-making. Through a large online experiment, we compare the utility of conformal prediction sets to displays of Top- and Top- predictions for AI-advised image labeling. In a pre-registered analysis, we find that the utility of prediction sets for accuracy varies with the difficulty of the task: while they result in accuracy on par with or less than Top- and Top- displays for easy images, prediction sets offer some advantage in assisting humans in labeling out-of-distribution (OOD) images in the setting that we studied, especially when the set size is small. Our results empirically pinpoint practical challenges of conformal prediction sets and provide implications on how to incorporate them for real-world decision-making.
@@ 1. Introduction
As machine learning models such as deep neural networks (NNs) achieve impressive predictive performance, AI-advised decision-making has become more commonplace in various situations where a human must make a decision, from high-stakes domains like medicine (e.g., (Razzak et al., 2018)) or autonomous vehicle navigation (e.g., (Fujiyoshi et al., 2019)) to online tasks like content moderation (e.g., (Lai et al., 2022)) or image labeling (e.g., (van der Wal et al., 2021)). Ideally, having access to a predictive model can help humans make more informed decisions, surpassing the capabilities of either human judgment or AI in isolation, a synergy known as complementary performance (Bansal et al., 2021; Hemmer et al., 2021). Whenever an AI system understands certain regions of the feature space better, the human counterpart can benefit by considering the model predictions. For example, in AI-advised image labeling, which we study in this paper, human labelers may benefit from viewing predictions from a model trained on previously labeled examples.
@@
To help humans know when to trust model predictions, presenting information about the probability that the model’s prediction is correct can, in principle, be useful. For example, a decision-maker could be provided with the softmax pseudo-probability produced in the final layer of NNs for classification tasks. However, NN predictions are known to be overconfident at times (Guo et al., 2017), and heuristic measures of uncertainty are often poorly calibrated. Consequently, decision-making relying on NN predictions with the highest softmax scores (e.g., Top- or Top-) can fail to account for prediction error, leading to worse decisions.
@@
Given these challenges, conformal prediction (Vovk et al., 2005; Romano et al., 2020; Angelopoulos et al., 2020) has emerged as an alternative solution that can rigorously quantify the prediction uncertainty for NNs through the use of prediction sets. Instead of presenting unreliable softmax values or other heuristics for uncertainty, the prediction set is a type of confidence interval consisting of a set of predictions with a coverage guarantee—the true class is, on average, captured within such sets with a user-specified probability (e.g., ). The conformal framework is compatible with NNs because the derived sets are distribution-free—they are model-agnostic, work with finite samples, and offer non-asymptotic guarantees without distributional assumptions (Angelopoulos and Bates, 2021).
@@
One challenge is that while conformal prediction sets express prediction uncertainty with guarantees that status quo presentations like Top- predictions with softmax cannot, the conformal coverage guarantee can lead to large sets embodying high uncertainty for instances the model perceives as difficult (Babbar et al., 2022; Angelopoulos et al., 2020). Compared with Top- predictions, such sets may increase cognitive load, rendering uncertainty quantification less effective (Zhou et al., 2017), thus compromising predictive performance. To understand how conformal prediction sets support human decisions aided by model predictions in an image labeling task, we contribute a large online repeated measure experiment () that compares the accuracy achieved in AI-advised image labeling using conformal prediction sets with that achieved with no predictions or Top- predictions. Participants in our study were tasked with labeling images from ILSVRC 2012 (Deng et al., 2009) that varied in difficulty and encompassed both image stimuli that were in-distribution and out-of-distribution (OOD). This allowed us to evaluate the utility of prediction sets against Top- presentation alternatives in scenarios where the model’s predictions were mostly accurate versus when the presence of “unknown unknowns” makes predictions more error-prone.
@@
We evaluate decision-making using metrics such as accuracy and the shortest path length, which approximates the amount of deviation from the ground truth in the label space hierarchy. We use participants’ elicited willingness-to-pay at the end of the experiment to compare the perceived value of predictions against the post-hoc inferred monetary benefits derived from prediction access.
@@
We find that for in-distribution instances, prediction sets lead to reduced labeling accuracy compared to Top- predictions. However, for OOD instances where model misspecification differentially affected the calibration of Top- predictions, prediction sets improve accuracy regardless of the set size. We find that participants are willing to pay roughly equivalent amounts for each type of display; if anything, they undervalue prediction sets relative to other presentations. Our results advance understanding of the strengths and limitations of prediction sets for improving AI-advised decision-making and highlight decision-makers’ tendency to over-rely on predictions, even when poorly calibrated, in an image labeling task.
@@ 2.1. Uncertainty Quantification Using Conformal Prediction
Prior work on Human-Computer Interaction (HCI) suggests that humans can make more informed decisions when uncertainty is effectively communicated (e.g., (Joslyn and LeClerc, 2012; Kay et al., 2016; Hullman et al., 2018)), such as by presenting prediction uncertainty through static intervals (Cumming, 2014; Cumming and Finch, 2005; Manski, 2019; Taylor and Kuyatt, 1994) or animations (Hullman et al., 2015; Zhang et al., 2021). However, uncertainty quantification for NNs remains challenging (Amodei et al., 2016; Angelopoulos and Bates, 2021) due to their black-box nature and unreliability of internal heuristic confidence values, such as softmax pseudo-probabilities (Guo et al., 2017; Torralba and Efros, 2011).
@@
One approach to quantify prediction uncertainty in classification tasks is using Bayesian NNs, where predictions depend on sampling from the posterior distributions of model parameters, carrying an inherent, mathematically grounded measure of uncertainty. However, Bayesian NNs heavily rely on distributional assumptions for their priors and can be computationally intensive to train. Conformal prediction (Vovk et al., 2005) has emerged as a more flexible and efficient solution that does not require strong model or distributional assumptions. In classification settings, conformal prediction quantifies uncertainty through prediction sets. Similar to traditional confidence intervals, the derived sets contain predicted classes and achieve a coverage guarantee that, on average, the true class is captured by the set with a user-specified probability. Formally, given a training set , where is the feature vector and is the corresponding label, the uncertainty set function, denoted , maps a feature vector to a subset of . This mapping ensures coverage111Recent work explores alternatives to marginal coverage, including class-conditional coverage (Gibbs et al., 2023) and relaxed versions of conditional coverage (e.g., (Hore and Barber, 2023)). that, for any new instance drawing from the same distribution as the training data, the probability over the randomness in the calibration and test points, with an error rate of (Angelopoulos and Bates, 2021).
@@
If an uncertainty quantification approach is well-calibrated, we naturally expect it to reflect greater uncertainty on more difficult instances for the model (e.g., OOD instances). With prediction sets (Romano et al., 2020; Angelopoulos et al., 2020), this manifests as adaptiveness: a conformal prediction set will be larger when the task is more difficult for the model. Adaptiveness is realized through the inductive conformal method, which quantifies the prediction uncertainty of a chosen classifier using a designated calibration set (Papadopoulos et al., 2002) separate from the training data set. This set contains instances that are independently and identically distributed (i.i.d.) to those on which the model is trained.
@@
For each calibration instance, the predicted softmax scores are adjusted using Platt scaling (Platt, 1999; Guo et al., 2017; Nixon et al., 2019) for better interpretability and further regularized to penalize unlikely classes (Angelopoulos et al., 2020). These calibrated probabilities are then ranked in descending order, reflecting the classifier’s confidence in each class. A conformity score is computed by summing these ranked probabilities up to and including the true class, capturing the model’s accumulated confidence for the instance leading up to the correct classification. From the distribution of the conformity scores of all calibration instances, a threshold is computed by taking the quantile after a finite sample correction for robustness. For a new and unknown instance, this calibration threshold decides the size of the prediction set by including all labels whose softmax scores exceed this calibration threshold. In our study, we opted for the Regularized Adaptive Prediction Sets (RAPS) algorithm (Angelopoulos et al., 2020) to ensure that the prediction sets of our experiment are as small as possible while achieving, on average, the desired coverage rate.
@@ 2.2. Effects of Presenting Uncertainty in AI-advised Decisions
Pairing humans with an AI system in complex decision-making scenarios can elevate performance beyond what either can achieve individually when their strengths are complementary (Bansal et al., 2021; Guo et al., 2024; Steyvers et al., 2022). In annotation tasks with large label spaces, AI models can improve accuracy and efficiency (Levy et al., 2021). Particularly, in AI-advised image labeling, various tools exist for applying automation to enhance human annotators’ speed and accuracy (Sager et al., 2021). However, performance on AI-advised tasks can be influenced by how predictions are presented—specifically, whether uncertainty is effectively communicated or not. Effective communication of uncertainty can improve the perceived trustworthiness of predictions (Zhang et al., 2020; Lai and Tan, 2019; Beller et al., 2013). Presenting a model’s accuracy is necessary to provide humans with a well-defined decision problem in many AI-advised decision settings (Hullman et al., 2024), and can encourage analytical thinking and reduce over-reliance on predictions (Zhang et al., 2020; Prabhudesai et al., 2023). However, uncertainty quantification does not guarantee improved decision-making. Cognitive biases, such as tendencies to under- or over-react to information relative to rational standards, play a major role in how uncertainty is interpreted (e.g., (Ambuehl and Li, 2018; Tversky and Kahneman, 1971)). Additionally, uncertainty quantification can improve trust under low cognitive load, but can diminish trust when uncertainty is presented ambiguously, demanding high cognitive load (Zhou et al., 2017).
@@
There is limited empirical evidence on how prediction sets influence trust and accuracy in AI-advised decision-making. One exception is Babbar et al. (2022), who investigated humans’ perceived utility and performance using RAPS versus Top- predictions for labeling images from CIFAR-100 (Krizhevsky, 2009). They found that participants using RAPS reported higher trust and perceived utility than those using Top- predictions. However, their study was unlikely to have achieved sufficient power to detect differences in accuracy between the two groups due to the very small sample size (15 participants per group). They suggest that smaller sets might lead to better accuracy, but this observation was based on a comparison between sets generated by two different conformal methods. Recent work by Straitouri and Rodriguez (2023) compared labeling performance between two decision support systems: one that only allowed participants to select a label value from the prediction set and another that allowed them to freely choose one of 16 labels. Their results suggest that when the model is well-calibrated, decision-making solely based on prediction sets tends to be more reliable and accurate because of the coverage guarantee.
@@
Motivated by overlapping research questions to Babbar et al. (2022), we conducted a much larger experiment to evaluate the utility of different uncertainty prediction displays for in-distribution and OOD stimuli, each systematically categorized by the difficulty and size of the associated prediction set. This enabled us to explore changes in performance by instance type. Contrary to the restricted agency in Straitouri and Rodriguez (2023), our study allowed participants to freely navigate a large label space, so as to allow for the possibility that even inaccurate predictions may serve as useful clues that help participants deduce the correct answer.
@@ 3.1. Overview
We conducted a large mixed-design repeated measure experiment on Prolific. Following a pre-registered222Pre-registration: https://aspredicted.org/6gh27.pdf analysis plan, we evaluated how different displays of prediction uncertainty impacted performance on an AI-advised labeling task, including relative to performance without access to predictions. Figure 1 provides an overview of our stimuli generation process, which systematically varied whether image instances were in-distribution or OOD and easy or difficult, as well as the size of conformal prediction sets.
This figure is an overview diagram illustrating the key experimental manipulations used in the study. It visualizes the process of utilizing conformal hold-out sets to create OOD images through synthetic image corruption, categorization of these images by difficulty and set size, and the subsequent selection of task images for the labeling tasks. The diagram further depicts the interface for different treatment conditions, including labeling without predictions (e.g., the baseline condition) or with access to various prediction displays with varying uncertainty quantification, across a series of tasks involving both in-distribution and OOD images.
@@
Each participant in our study labeled a set of 16 images sampled from the ILSVRC 2012. We assigned participants to one of four conditions representing different prediction displays: (1) no predictions (baseline), (2) Top- prediction with softmax, (3) Top- predictions with softmax, or (4) the RAPS prediction set (Angelopoulos et al., 2020).
@@
In real-world scenarios, the deployed model can encounter unexpected OOD instances where the distribution of input features differs from what was trained and calibrated. We varied whether images were in-distribution or OOD within subjects. Among the 16 stimuli, 4 were in distribution, while 12 were not, as we expected greater variation in participants’ accuracy for OOD.
@@
In our experiment, OOD instances notably lowered the model’s Top- prediction accuracy. Although the corrupted images based on the ILSVRC 2012 dataset that we used as OOD instances could compromise the conformal coverage guarantees, the adaptiveness of RAPS led to good coverage despite OOD. However, the number of labels included in the prediction sets became much larger when the instance was harder for the model. To evaluate the impact of set size on the effectiveness of prediction sets, we balanced set size within subject by categorizing images as smaller or larger based on the size of the prediction set generated for the image. This factor enabled us to examine the performance of the prediction set between size levels, as previous work suggests that a smaller set size is more preferable (Romano et al., 2020; Angelopoulos et al., 2020) and may yield higher accuracy for in-distribution images (Babbar et al., 2022). We expected smaller set sizes to be more useful in general, as they could reduce cognitive load while maintaining the conformal coverage guarantee for in-distribution images. Larger sets embodying high uncertainty could still be useful for labeling when participants were highly uncertain of what the image showed and consequently heavily relied on prediction sets. In our setting, set size could also be viewed as a nuanced measure of difficulty for in-distribution stimuli because images that required a larger set size to achieve the desired coverage were intrinsically more difficult (Angelopoulos et al., 2020).
@@
For both in-distribution and OOD stimuli, we balanced difficulty within subject by easy and hard instances, using the median cross-entropy loss of in-distribution images as a threshold to define these levels. The former enabled us to assess whether participants were inclined to accept predictions that expressed high confidence. The latter allowed us to examine whether participants were more likely to modify or reject low-confidence predictions and instead rely on their own intuition.
@@
Our study evaluated participants’ performance by their labeling accuracy and the shortest path length between the chosen label and the correct label in the WordNet hierarchy (Miller, 1995; Fellbaum, 1998). After they had labeled all images, we asked them to report their willingness-to-pay for the prediction display they used (i.e., a value ranging from to USD, the maximum they could earn) if they were to repeat the study with a similar distribution of instances. Finally, we elicited the strategies employed to approach the tasks.